OpenStudy (anonymous):

An object with a mass of 1 kg is dropped from rest from a height of 4 meters above the ground. Its total mechanical energy is _____________. Show your work and units! HALP PLS ASAP I`LL DO ANYTHING

OpenStudy (anonymous):

Mechanical energy=PE +KE =mgh +1/2mv^2 =1×10×4+0 =40 joules